# LOGen: Toward LiDAR Object Generation by Point Diffusion

### v2

@@ 4 Experiments
4.1 Experimental Setup
Dataset.
We evaluate our approach on nuScenes [3], which was acquired with a 32-beam LiDAR, thus creating notably sparse sweep patterns. To construct object samples, we use the semantic point labels and the 3D box annotations of objects. We take as object instance all points falling into a box, filtering out possible points of a different class. We also discard objects with less than 20 points. Point intensity, which ranges in 0–255, has a highly irregular distribution, with most examples clustered from 0–30, and with outliers. To create a more uniform intensity distribution, we use log max scaling.
In total, we extract 268,684 objects with at least 20 points which covers more than 95% of all object points. It can also be noted that the official instance segmentation evaluation of nuScenes only considers instances with at least 15 points [3]. In order to unify the number of points per cloud in a batch, we pad each point cloud to the max length in the batch. For all experiments we use the conditioning information from objects in the validation set, approximating novel views of objects.
Object-level metrics.
To evaluate the ability to generate a sensible variant of a real object given only the conditioning information corresponding to the real object, we use two pointset metrics: Chamfer Distance (CD) and Earth Mover’s Distance (EMD). These metrics penalize incoherent shape generations which diverge from original objects, however also penalizing realistic but diverse generations. To then evaluate the realism of generations, we use different distribution-based metrics which compare the set of real objects of a given class to a corresponding set of objects generated with the associated conditioning.
We use the Fréchet PointNet Distance (FPD) [37], an adaptation of the classic Fréchet Inception Distance (FID) [12] by replacing 2D Inception features with 3D PointNet features [33]. However, the original FPD used as reference relies on the original PointNet classifier, trained on ModelNet40 [33]. While it is well suited for use on objects generated from a 3D shape dataset, with dense and uniform point sampling, it is not for LiDAR-scanned objects. We thus use instead a separate PointNet classifier trained on the real LiDAR objects of nuScenes.
We also measure the Accuracy of that same PointNet Classifier (APC), i.e., trained on real objects, for recognizing synthetic objects. It is to be compared to the accuracy of the classifier for recognizing real objects.
We additionally consider the Kernel Inception Distance (KID), which has advantages over FID, in particular for small datasets [2]. Like for FID, we adapt KID into a Kernel PointNet Distance (KPD), replacing the Inception network by our trained PointNet.
Besides the above metrics, which compare the distributions of real and synthetic objects at feature level, we also consider distribution comparisons at a pointset level. We use coverage (COV) [1], which measures the proportion of real pointsets matched to at least one generated pointset, and 1-nearest neighbor accuracy (1-NNA) [52], derived from [23], which merges real and synthetic objects and counts the proportion of closest neighbors of each sample that are of the same source. 1-NNA ranges from 100% when distributions are separate, to 50% when they are indistinguishable. It evaluates both diversity and quality in a single metric. COV and 1-NNA are measured with CD and EMD.
We extend COV and 1-NNA to measure the quality of the generated intensities. We extract 256-dimensional normalized intensity histograms from generated objects to use as features for COV and 1-NNA computations. The Appendix details all metrics.
Model 1-NNA(%) $\downarrow$ COV(%) $\uparrow$ Intensities
[…]
MinkUNet - 0.208 0.212 79.5 74.4 33.7 41.7 N/A N/A
PixArt-L PE 4 ch 0.196 0.204 76.5 72.5 33.5 41.8 71.6 32.9
DiT-3DL VE 3 ch 0.136 0.121 75.8 70.9 34.3 42.0 N/A N/A
DiT-3DL PE 4 ch 0.142 0.125 73.4 70.7 35.5 42.1 74.4 31.4
LOGen PE 4 ch 0.130 0.111 74.1 70.6 36.0 42.3 68.2 37.0
Table 1: Point-based metrics. Mean classwise score for CD, EMD, 1-NNA, COV. Emb.: type of point embedding: none (-), voxel-based (VE 3 ch) or PointNet-based (PE 4 ch). The last two columns report 1-NNA and COV scores for intensity histograms based on CD.
Scene-level metrics.
A second set of evaluations is conducted on the perception of generated objects in a scene context by a typical 3D segmentation model with the contention that more realistic objects should be segmented at a similar rate to real objets. We consider the replacement of ground-truth objects by objects generated with the same conditioning, then measure the change in classwise mIoU segmentation performance of the model on this modified data. We use a variant of nuScenes, nuScenes (LOGen), where each object with at least 20 points is replaced by a generated object of the same class and box size, seen under the same viewing angle, and with the same number of points.
Model FPD $\downarrow$ KPD (%) $\downarrow$ APC (%) $\uparrow$
[…]
MinkUNet - 25.4 N/A 3.18 N/A 22.6 N/A
DiT-3DL VE 3 ch 2.80 N/A 0.25 N/A 39.8 N/A
DiT-3DL PE 4 ch 2.62 6.99 0.25 0.75 37.2 41.7
PixArt-L PE 4 ch 2.95 5.16 0.24 0.41 37.8 45.5
LOGen PE 4 ch 1.34 2.18 0.10 0.12 40.9 48.1
Real dataset objects - - - - 39.0 46.4
Table 2: Feature-based metrics. Mean classwise scores for FPD, KPD, APC. Emb.: type of point embedding: none (-), voxel-based (VE 3 ch) or PointNet-based (PE 4 ch).
Distance FPD $\downarrow$ KPD (%) $\downarrow$ APC (%) $\uparrow$
to sensor 3 ch. 4 ch. 3 ch. 4 ch. 3 ch. 4 ch.
orig. ($d$) 1.34 2.18 0.10 0.12 40.9 48.1
$d\times 2$ 4.79 12.70 0.61 1.71 34.0 38.4
$d\mathbin{/}2$ 3.08 8.13 0.33 0.85 38.3 49.8
Table 3: Evaluation of conditioning on distance to sensor. LOGen generates realistic instances at varying distances. Note that the distribution of objects has diverged from the GT as objects are more or less dense according to the point size scaling, especially in the $d\times 2$ case, where the point clouds have become much sparser.
Baselines.
Given the novelty of the task, we construct the following baselines: MinkUNet is the sparse convolutional architecture used in LiDiff [31], which ignores intensities. PixArt-L is our adaptation of PixArt-$\alpha$ to LiDAR points and conditioning, using our PointNet embeddings (PE) to allow generation of intensities. DiT-3DL VE 3 ch is our adaptation of DiT-3D to LiDAR conditioning, i.e., with its original voxel embeddings (VE) and no handling of intensities. (The voxelization library used in DiT-3D [27] does not handle a fourth input channel; handling intensities would require adapting the voxelization layer like in MinkUNet [6], as separate input features.) DiT-3DL PE 4 ch replaces VE by our PE in DiT-3DL VE 3 ch, and thus also handles intensities.
Model Emb. bar. bic. bus car c.veh mot. ped. t.con. tra. tru. mean
3 channel MinkUNet 76.1 31.4 18.7 84.4 14.7 63.6 69.8 59.3 12.4 49.7 48.0
DiT-3DL VE 3 ch 48.4 42.5 0.3 34.1 0.2 63.8 75.6 69.6 7.7 3.4 34.6
DiT-3DL PE 4 ch 61.6 39.2 11.1 84.2 1.7 68.8 77.5 51.3 15.9 26.6 43.9
PixArt-L PE 4 ch 58.9 32.1 3.9 75.7 0.9 76.6 79.0 56.7 13.6 51.3 44.9
LOGen PE 4 ch 78.2 50.9 60.1 79.1 28.1 82.2 84.2 73.5 37.5 63.2 63.7
GT 80.3 48.8 90.7 94.2 40.6 86.0 85.5 73.4 60.7 85.2 74.5
PixArt-L PE 4 ch 56.6 33.7 14.5 82.4 8.1 78.5 79.3 62.8 26.7 52.6 49.5
LOGen PE 4 ch 78.0 48.9 71.7 91.3 37.1 82.0 84.8 79.1 45.1 72.9 69.1
GT 81.6 50.4 90.9 94.6 44.5 88.1 85.6 77.4 64.6 85.6 76.3
Table 4: Scene-level comparison. SPVCNN is trained with 3 or 4 channels (coordinates + intensity) on nuScenes (train set) with real objects. It is tested on nuScenes (val set) where each object ($\geq$ 20 points) is replaced with a generated object of the same class, box size and sensor viewing angle, and with the same number of points, and where objects with less than 20 points are removed from the dataset. Also note that, intensities generated by LOGen are realistic enough to provide +5.4 mIoUat scene level.
4.2 Object-level evaluation
Qualitative generation examples are shown in Fig. 3 and Fig. 4.
Point embeddings.
Tab. 1 and 2 show that the voxel embeddings of DiT-3D do not perform as well as our PointNet embeddings, which supports our proposal (Sec. 3.3). MinkUNet succeeds in generating local structure (LiDAR pattern), but fails to generate objects respecting size conditions. The lack of point embeddings in MinkUNet create global size inconsistencies.
Architecture.
Tab. 1 shows that LOGen transformer blocks perform slightly better than the blocks proposed in PixArt-$\alpha$ and DiT-3D. Tab. 2 shows that the superiority of LOGen blocks is more pronounced as measured by the feature-based metrics of FPD and KPD.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Examples of generated intensities. Clockwise: Truck, Bus, Motorcycle, Traffic Cone. For each pair, the left image is from the nuScenes validation set, the right image is generated using the same conditioning information and number of points. More examples can be seen in Appendix.
Evaluation of conditioning on distance to sensor.
For all dataset statistics, class-wise results, and a comparison of single-model-per-class vs. multi-class approaches, please refer to the Appendix.
4.3 Scene-level evaluation
Semantic quality of generated data.
Tab. 4 shows that the LOGen generations are more semantically identifiable from the point of view of an SPVCNN segmenter. Though trained on real objects, the segmenter is able to recognize LOGen objects at a much higher mIoU than other models, +19.6 points versus the second best model. This is particularly pronounced for rare classes such as bicycle, bus, trailer, and construction vehicle.
Test Box mIoU (%) $\uparrow$
objects orientation 3 ch. 4 ch.
real orig. 74.5 76.3
LOGen orig. 63.7 69.1
LOGen rot. $\nicefrac{{1}}{{5}}$ 53.7 64.6
LOGen rot. $\nicefrac{{2}}{{5}}$ 54.8 66.0
LOGen rot. $\nicefrac{{3}}{{5}}$ 53.8 63.8
LOGen rot. $\nicefrac{{4}}{{5}}$ 53.3 64.1
Table 5: Evaluation of conditioning on sensor viewing angle. SPVCNN is trained with 3 or 4 channels (coordinates + intensity) on nuScenes with real objects. It is tested on nuScenes where each object ($\geq$ 20 points) is replaced with a LOGen-generated object of the same class, box size, and sensor viewing angle, but rotated on itself.
Evaluation of conditioning on sensor viewing angle.
To assess LOGen’s ability to generate realistic objects regardless of viewing angle, we modify the synthetic replacements by rotating their angular condition. This reduces the need to find suitable locations for new objects. We consider five evenly spaced angles around a full turn, i.e., $\nicefrac{{i}}{{5}}\times 360$° where $i\in{0..4}$, as shown in Fig. 1.
Tab. 5 shows a moderate variation of the performance when rotating the objects. This performance is partly due to collisions between objects that occur after rotation (the transformation is performed without such a test). Further, the distribution of object viewing angles in the dataset is not uniform. Less viewed angles can result in noisier generations.
Limitations.
Currently we train from occluded objects with partly empty bounding boxes without any modifications. Consequently, all objects we generate tend to fill the conditioned bounding box, and thus fill in areas of objects that should be occluded.
@@ 5 Conclusion
This work demonstrates that a well-designed diffusion model is capable of generating realistic LiDAR objects. Novel metrics, baselines and LOGen (best performing architecture) establish the first standard in LiDAR object generation. Future work can increase generation speed and memory efficiency of the models, and develop novel methods to use generated objects.
@@ A.3 Denoising neural network
[…]
@@
$p(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\bm{\mu}_{q}(x_{t},x_{0}),\mathbf{\Sigma}_{q}(t))$ (2)
where $\bm{\mu}_{q}$ is a function of both $x_{t}$ and $x_{0}$, and $\mathbf{\Sigma}_{q}(t)=\sigma_{q}^{2}(t)\mathbf{I}$ where
@@ A.4 Guiding the diffusion process
Under the continuous-time interpretation, the diffusion process can be described as a direction through the latent space, from a standard Gaussian Distribution towards the data generating distribution [39]. This direction formulation ties directly into the use of conditioning information to guide the diffusion process. The conditioning information changes the direction through the latent spaces towards a specific part of the data generating distribution.
@@
Ultimately it is not necessary to train two different diffusion models. During the training procedure, the conditioning information is replaced by a null token in a manner which imitates random dropout. A single diffusion model is thus capable of performing Classifier Free Guidance [24].
@@ A.5 Diffusion transformer
[…]
Figure 7: Architecture of LOGen and baselines PixArt-L and DiT-3DL.
@@
Our baselines DiT-3DL and PixArt-L, whose architecture is shown Fig. 7 together with LOGen, are directly adapted from their original counterparts to handle: (i) LiDAR data, i.e., point embedding, and (ii) LiDAR object conditioning. LOGen is then a modification of PixArt-L, as described in main paper.
@@ Appendix B Implementation details
[…]
Training. The Adam optimizer was used with a constant learning rate of $0.0001$. We trained small classes (i.e., barrier, bike, car, motorcycle, pedestrian, traffic cone) for 500,000 iterations with a batch size of 16. The remaining large object classes (i.e., bus, construction vehicle, trailer and truck) are trained for 5,000,000 iterations with a batch size 4. All trainings were conducted on 4 Nvidia A100 GPUs.
@@ Appendix C Metrics
  • Chamfer distance (CD) and Earth mover distance (EMD):
@@
    $\textrm{CD}(X,Y)=$
    $\quad\sum_{x\in X}\min_{y\in Y}||x-y||^{2}_{2}+\sum_{y\in Y}\min_{x\in X}||x-y||^{2}_{2}\>,$ (8)
    $\text{EMD}(X,Y)=\min_{\gamma:X\rightarrow Y}\sum_{x\in X}||x-\gamma(x)||_{2}\>,$ (9)
@@
  • 1-nearest neighbor accuracy (1-NNA):
@@
    $\quad\frac{\sum_{X\in S_{g}}\mathbb{I}[N_{X}\in S_{g}]+\sum_{Y\in S_{r}}\mathbb{I}[N_{Y}\in S_{r}]}{|S_{g}|+|S_{r}|}\>,$
  • Fréchet Point Cloud Distance (FPD)
    FPD [37] is an adaptation of the classic Fréchet Inception Distance (FID) [12], where 2D Inception features are replaced with 3D PointNet features [33]. Like FID, FPD computes the 2-Wasserstein distance between the real and generated Gaussian distributions in the feature space.
    $\text{FPD}(S_{r},S_{g})=$
    $\quad||\mathbf{m}_{S_{r}}-\mathbf{m}_{S_{g}}||_{2}^{2}+\text{Tr}(\Sigma_{S_{r}}\,{+}\,\Sigma_{S_{g}}\,{-}\,2(\Sigma_{S_{r}}\Sigma_{S_{g}})^{\frac{1}{2}}),$ (12)
@@
    The original FPD is based on the PointNet classifier trained on ModelNet-40 [33] with dense and uniform point sampling. To account for the sparse and non-uniform characteristics of LiDAR-scanned objects in nuScenes, we instead use a separate PointNet classifier that has been trained on the real LiDAR data from nuScenes. Like with the original FPD, we extract a 1803-dimensional feature vector (5 fewer dimensions than the original 1808, due to the presence of 5 fewer classes) from the output of the dense layers, to calculate the mean and covariance in Appendix C.
  • Kernel Point Cloud Distance (KPD)
    We define KPD by adapting the Kernel Inception Distance (KID) [2], where 2D Inception features are replaced with our 1803-dimensional 3D PointNet features. Like KID, KPD calculates the maximum mean discrepancy (MMD) between the real and generated instances in the feature space.
    $\text{KID}(S_{r},S_{g})=\text{MMD}(S_{r},S_{g})^{2}\>.$ (13)

### v3

@@ 4.3 Scene-level evaluation
barrierp
bikep
busp
carp
constr. veh.p
motorcyclep
pedestrianp
traffic conep
trailerp
truckp
Mean barrier bicycle bus car constr. veh. motorcycle pedestrian traf. cone trailer truck
$\textbf{0.3}$ 0.5-0.5 0.8-0.8 1.2-1.2 +0.7+0.7 1.7-1.7 +0.5+0.5 0 +4.3+4.3 +0.3+0.3 +1.1+1.1
Table 4: Impact of inplace data augmentation with LOGen samples for semantic segmentation on nuScenes validation set. During training, each sample is selected with a 50% probability to be either a real sample or a generated one. We display the class-wise increase in IoU%.
4.4 Downstream evaluation
Object-level augmentation for semantic segmentation on nuScenes.
To further motivate the focus on object level generation, we design a simple experiment highlighting one use case of the generated objects, enhancing diversity of objects in the training dataset. We train the SPVCNN [44] model for semantic segmentation on the nuScenes dataset. During training, with a 50% probability, we replace real objects with versions generated by LOGen. Results are in Tab. 4. We highlight the improvements for specific rare classes such as traffic cones, trailers, and trucks, and further improvements for cars. Note that, in this downstream task, using a single-class model is consistent with the selection of specific rare classes to augment.
Cross-dataset object augmentation.
We use LOGen trained on nuScenes to enhance the semantic segmentation performance on SemanticKITTI for classes pedestrian and motorcycle, which are commonly targeted for augmentation due to the relatively low performance of their segmentation. We randomly insert 10 such generated objects for each class into SemanticKITTI scenes during SPVCNN training. This simple augmentation strategy improves classwise IoU% from 64.8 to 67.3 for motorcycle and from 65.0 to 66.9 for pedestrian.
We provide a comprehensive supplementary material that includes detailed information on the diffusion background, implementation specifics, official metric definitions, dataset statistics, evaluation of conditioning parameters, all classwise performance results, comparisons between LOGen’s single-model-per-class and multi-class approaches, qualitative comparisons between LOGen and baselines, examples of novel LOGen objects and videos showcasing our object generation results.
4.5 A discussion on scalability
Following other works in the object generation literature [30], we mainly chose to train and evaluate a single model per class. This allowed us to use a compact network, and avoid inconsistent batch sizes. Due to the use of zero-padding, mixing objects with high variation in size results in certain objects being padded with many zero elements. We note that training single model per class introduces no drawbacks in terms of runtime or model efficiency when evaluated in iterations per instance.
We train and evaluate a multi-class model of the same size trained with the same iterations as shown in Table 3, where the multi-class model does not perform as well as the single class versions under most perceptual metrics. However the multi-class model does have the best overall 1-NNA and Coverage, indicating avenues for future work.
Further, while the point embeddings outperform the voxelization approaches in terms of generated quality, the lack of any point compression creates memory consumption issues for larger objects.
@@ 5 Conclusion
This work demonstrates that a well-designed diffusion model is capable of generating realistic LiDAR objects. Novel metrics, baselines and LOGen (best performing architecture) establish the first standard in LiDAR object generation. Interesting perspectives include further development of downstream applications. Future work can increase generation speed and memory efficiency of the models, and develop novel methods to use generated objects.
Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocations AD011013503R1 and AD011013503R3 made by GENCI.
Appendix
This appendix is organized as follows.
  • Appendix A provides some background on diffusion and in particular on transformer-based diffusion.
  • Appendix B provides more implementation details.
  • Appendix C offers definitions and formulas for the metrics we used.
  • Appendix D provides various statistics regarding the distribution of objects we evaluate on.
  • Appendix E evaluates the impact of interpolating the conditioning information to produce novel objects.
  • Appendix F gives class-wise results.
We also provide videos corresponding to generated examples from each class on this webpage https://nerminsamet.github.io/logen.
@@ A.3 Denoising neural network
We define the forward destructive process with a distribution $q(x_{t}|x_{t-1})$ at step $t$ for a dataset sample $x$. The associated Markov chain $\{x_{0}\,...\,x_{t}\,...\,x_{T}\}$ is defined such that $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$ where $0\,{<}\,...\,{<}\,\beta_{t}\,{<}\,...\,\beta_{T}\,{<}\,1$ defines a scaling schedule for the applied noise. This forward process can be re-written such that each time step depends only on the original data via the introduction of $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ such that $q(x_{t}|x_{o})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I}))$. Given Gaussian noise $\epsilon_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the reparameterization trick can then be recursively applied to derive a closed form of any $x_{t}$ as explained in [27]:
@@
$p(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{q}(x_{t},x_{0}),\Sigma_{q}(t))$ (2)
where $\bm{\mu}_{q}$ is a function of both $x_{t}$ and $x_{0}$, and $\Sigma_{q}(t)=\sigma_{q}^{2}(t)\mathbf{I}$ where
@@ A.4 Guiding the diffusion process
Under the continuous-time interpretation, the diffusion process can be described as a direction through the latent space, from a standard Gaussian Distribution towards the data generating distribution [42]. This direction formulation ties directly into the use of conditioning information to guide the diffusion process, changing the direction through the latent spaces towards a specific part of the data generating distribution.
@@
Ultimately it is not necessary to train two different diffusion models. During training, the conditioning information is replaced by a null token in a manner which imitates random dropout. A single diffusion model is thus capable of performing Classifier Free Guidance [27].
@@ A.5 Diffusion transformer
The Diffusion Transformer has been adapted for a 3D shape generation task, where it achieves state-of-the-art results [30]. This DiT-3D introduce a voxelization step, which converts the unstructured 3D point cloud data into features of a $\times\,32$$\times\,$32 voxel grid. These voxels are then grouped into patches and embedded. This makes it possible to use more or less the same overall architecture as the Diffusion Transformer, using the same adaptive Layer Norm layers to apply conditioning information (cf. Fig. 5 (right) of this appendix). This allows the leveraging of models pre-trained on large 2D datasets for 3D generative tasks.
@@
Our baselines DiT-3DL and PixArt-L, whose architecture is shown Figure 2 (of main paper) together with LOGen, are directly adapted from their original counterparts to handle: (i) LiDAR data, i.e., point embedding, and (ii) LiDAR object conditioning. LOGen is then a modification of PixArt-L, as described in main paper.
@@ Appendix B Implementation details
Diffusion parameters.
The DDPM scheduler was used with 1000 time steps during training. During inference 500 time steps were used. The $\beta$ schedule was linearly spread on the range of $3.5\ 10^{-5},0.007]$, as in LiDiff. The parameter $\lambda$ controlling the strength of the guidance was set to 1.0 (best value from a grid search, while LiDiff uses $\lambda=6.0$). This value of $\lambda$ cancels out the unconditional score term in Eq. 7. However, this equation is only used during the sampling procedure (which we do at inference or validation). During training, we periodically use null conditions for 10% of the conditions in a batch (as LiDiff does). So the classifier free guidance does have an impact on the training procedure.
Training.
The Adam optimizer was used with a constant learning rate of $0.0001$. We trained small classes (i.e., barrier, bike, car, motorcycle, pedestrian, traffic cone) for 500,000 iterations with a batch size of 16. The remaining large object classes (i.e., bus, construction vehicle, trailer and truck) are trained for 5,000,000 iterations with a batch size 4. All trainings were conducted on 4 Nvidia A100 GPUs.
@@ Appendix C Metrics
Chamfer distance (CD), Earth mover distance (EMD).
@@
[…]
@@
1-nearest neighbor accuracy (1-NNA).
@@
$\hskip-10.243pt\textrm{1-NNA}(S_{g},S_{r})=\\ \frac{\sum_{X\in S_{g}}\mathbb{I}[N_{X}\in S_{g}]+\sum_{Y\in S_{r}}\mathbb{I}[N_{Y}\in S_{r}]}{|S_{g}|+|S_{r}|}\>,$ (11)
Fréchet Point Cloud Distance (FPD).
FPD [40] is an adaptation of the classic Fréchet Inception Distance (FID) [14], where 2D Inception features are replaced with 3D PointNet features [36]. Like FID, FPD computes the 2-Wasserstein distance between the real and generated Gaussian distributions in the feature space:
@@
The original FPD is based on the PointNet classifier trained on ModelNet-40 [36] with dense and uniform point sampling. To account for the sparse and non-uniform characteristics of LiDAR-scanned objects in nuScenes, we instead use a separate PointNet classifier that has been trained on the real LiDAR data from nuScenes. Like with the original FPD, we extract a 1803-dimensional feature vector (5 fewer dimensions than the original 1808, due to the presence of 5 fewer classes) from the output of the dense layers, to calculate the mean and covariance in Eq. 12.
Kernel Point Cloud Distance (KPD).
Likewise, we adapt the Kernel Inception Distance (KID) [3], replacing 2D Inception features with our 1803-dimensional 3D PointNet features. Like KID, KPD calculates the maximum mean discrepancy (MMD) between the real and generated instances in the feature space:
$\text{KPD}(S_{r},S_{g})=\text{MMD}(S_{r},S_{g})^{2}\>.$ (13)
Accuracy of PointNet Classifier (APC).
We also measure the accuracy of the PointNet classifier, i.e., trained on real objects, for recognizing synthetic objects. It is to be compared to the accuracy of the classifier for recognizing real objects.